HOW MANY GRAMS SiO2 IS REQUIRED TO PRODUCE 73.8G OF SiC ?


1
Expert's answer
2022-02-21T17:13:06-0500

Solution:

Calculate the moles of SiC:

The molar mass of SiC is 40 g/mol

Therefore,

Moles of SiC = (73.8 g SiC) × (1 mol SiC / 40 g SiC) = 1.845 mol


Balanced chemical equation:

SiO2 + 3C → SiC + 2CO

According to stoichiometry:

1 mol of SiO2 produces 1 mol of SiC

X mol of SiO2 produces 1.845 mol of SiC

Thus,

Moles of SiO2 = X = (1.845 mol SiC) × (1 mol SiO2 / 1 mol SiC) = 1.845 mol SiO2


Calculate the mass of SiO2:

The molar mass of SiO2 is 60 g/mol

Therefore,

Mass of SiO2 = (1.845 mol SiO2) × (60 g SiO2 / 1 mol SiO2) = 110.7 g SiO2

Mass of SiO2 = 110.7 g


Answer: 110.7 grams of SiO2 is required.